OpenStudy (anna_kath_6):

he zero product property says that if ab = 0 then either a or b or both equal zero. So here when (x - 4) or (x - 3) equal zero you have your x intercept

jhonyy9 (jhonyy9):

for example for x-4=0 and to both sides 4 and what will get ?

jhonyy9 (jhonyy9):

sorry add to both sides 4

jhonyy9 (jhonyy9):

x-4+4 = 0+4 do you can continue it now ?

jhonyy9 (jhonyy9):

bc. -4+4=0 x=+4 x=4 right ?

jhonyy9 (jhonyy9):

do you can solve it now on the second case ?

jhonyy9 (jhonyy9):

x-3=0 add to both sides +3 so what will get ?

jhonyy9 (jhonyy9):

courage please this is easy exactly same with above wrote in case of x-4=0 just here not is 4 is 3